CONTINUE THIS NUMBER PATTERN:
1,8,27,64,__, __, __,__, __,1000...

Answer by amoati(11)

1,8,27,64,__, __, __,__, __,1000
8 = 2 * 2 * 2
27 = 3 * 3* 3
64 = 4 * 4 *4
so the next number would be 5 * 5 * 5 = 125
and so now so the series would be
1,8,27,64,125,216,343,512,729,1000